6.12 L03 AN) Dalton Inc. produces and sells three products Unit data concerning each product is shown below Product $200 Selling price Direct labor costs Other variable costs 30 95 $300 80 80 $250 35 145 The company has 2,000 hours of labor available to build inventory in anticipation of the companys peak season. Management is trying to decide which product shouldbe produced. The direct labor hourly rate is S10. Instructions a. Determine the number of direct labor hours per unit. b. Determine the contribution margin per direct labor hour. c. Determine which product should be produced and the total contribution margin for that product.
Answer #1

Answers

  • All working forms part of the answer
  • Requirement ‘a’

Working

D

E

F

A

Direct Labor Cost

$                   30.00

$                 80.00

$                 35.00

B

Direct labor hourly rate

$                   10.00

$                 10.00

$                 10.00

C = A/B

No. of direct labor hours per unit

3

8

3.5

  • Requirement ‘b’

Working

D

E

F

A

Selling price

$                200.00

$               300.00

$               250.00

B

Direct labor cost

$                   30.00

$                 80.00

$                 35.00

C

Other variable costs

$                   95.00

$                 80.00

$               145.00

D = A - B - C

Contribution margin per unit

$                   75.00

$               140.00

$                 70.00

E [calculated above]

No. of direct labor hours per unit

3

8

3.5

F = D/E

Contribution margin per direct labor hour

$                   25.00

$                 17.50

$                 20.00

  • Requirement ‘c’

Answer = Product D should be produced because it gives maximum ‘contribution margin per direct labor hours’ of $ 25

Total Contribution margin for that product = 2000 hours x $ 25 = $ 50,000